OpenStudy (anonymous):

sin^2(x)cos^2(x)-cos^2(x)

OpenStudy (anonymous):

hard to tell exactly what you want done here, but i would start be factoring out a cos^2(x)

OpenStudy (anonymous):

\[\cos^2\theta(\sin^2\theta-1)\]\[\cos^2\theta(-\cos^2\theta)=-\cos^4\theta\]?

OpenStudy (anonymous):

What are you trying to do?

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

Yes that is correct.

OpenStudy (anonymous):

Sin^2 + cos^2 = 1 Sin^2 - 1 = cos^2. Just had to use this on my calculus III test :(